# Methods Paper Version 19

### v1

@@ IV Two-Channel Scattering
We have shown that the Efros method is consistent and leads to a convergence of single-channel scattering calculations using a limited basis space obtained by diagonalizing the truncated Hamiltonian $H^{tr}$. We now extend our application to a system with two open channels. In particular, we investigate $np$ scattering in the ${}^{3}SD_{1}$ channel using the Daejeon16 $NN$ potential and eigenfunction SRFs.
IV.1 Hamiltonian, $S$ matrix and Stapp parameterization for $np$ scattering in coupled partial waves
@@
$H={\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle V_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|+{\color[rgb]{0,0,0}\sum_{i=1,2}}\>\sum_{n,n^{\prime}=0}^{\infty}|\phi_{ni}\rangle T^{l_{i}}_{nn^{\prime}}\langle\phi_{n^{\prime}i}|\equiv{\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle H^{tr}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|\\ +\sum_{i=1,2}\>\sum_{n,n^{\prime}=\mathcal{\color[rgb]{0,0,0}N}_{i}}^{\infty}|\phi_{ni\!}\rangle T^{l_{i}}_{nn^{\prime}}\langle\phi_{n^{\prime}i}\!|\!+\sum_{i=1,2}\!\big{[}|\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1,i\!}\rangle T^{l_{i}}_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1,\mathcal{\color[rgb]{0,0,0}N}_{i}}\!\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}i}\!|+|\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}i}\rangle T^{l_{i}}_{\mathcal{\color[rgb]{0,0,0}N}_{i},\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1,i}|\big{]}\!.$ (72)
We define the truncation of the potential energy matrix by the maximal oscillator quanta $N_{\max}$, i. e., ${2n+l_{i}}\leq N_{\max}$ and $\mathcal{\color[rgb]{0,0,0}N}_{i}=\frac{1}{2}(N_{\max}-l_{i})+1$, $i=1,2$.
@@
$\tilde{V}\equiv\sum_{i,j=1,2}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle\tilde{V}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|=\sum_{i,j=1,2}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle\sigma_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}^{n}V_{ni,n^{\prime}j}\sigma_{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}^{n^{\prime}}\langle\phi_{n^{\prime}j}|.$ (73)
Smoothing also improves the convergence in the two-channel case. We are not presenting results prior to smoothing but only present below results of calculations with the smoothed interactions. That means that we use the truncated Hamiltonian
$H^{tr}\equiv\tilde{H}^{tr}={\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle H^{tr}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|$ (74)
@@
For the two-channel problem, it is convenient to solve Eq. (22) in the $S$-matrix representation. We define the set of SRFs $\{\beta_{q}\}$ as a set of $N-2$ eigenfunctions of $H^{tr}\!$,
@@
Note, $N=v+2$ is the number of unknowns in each channel and the total number of independent unknowns is ${{\mathbb{\color[rgb]{0,0,0}N}}=2N-1\color[rgb]{0,0,0}=2v+3}$. We first solve the full set of equations for the $i=1$ entrance channel with the set $\{\bar{\beta}_{q}^{1}\}$ constituting a set of $N-2$ eigenfunctions of $H^{tr}$ plus two additional oscillator functions that are used exclusively for orthogonalization and guarantee coupling to the continuum but do not appear explicitly in Eq. (1):
${\color[rgb]{0,0,0}\bar{\beta}_{q}^{1}(\vec{r})}=\begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...,N-3,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{1}1}(\vec{r}),&q=N-2,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{2}2}(\vec{r}),&q=N-1.\end{cases}$ (77)
@@
$\beta_{q}^{\textrm{Efn}}(\vec{r})=\sum_{q^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{1}-1}a_{qq^{\prime}}^{l_{1}}\,{\color[rgb]{0,0,0}\phi_{q^{\prime}1}(\vec{r})}+\sum_{q^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{2}-1}a_{qq^{\prime}}^{l_{2}}\,\color[rgb]{0,0,0}\phi_{q^{\prime}2}(\vec{r}).$ (78)
Next, we solve the reduced set of equations for the ${i=2}$ entrance channel. Since there is one less unknown, namely, the off-diagonal $S$-matrix element ${\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)}$, we use for the set $\{\bar{\beta}_{q}^{2}\}$ a subset of $\{\bar{\beta}_{q}^{1}\}$ that includes one less function. The choice of the function to be omitted for the orthogonalization in the $i=2$ channel is arbitrary. We can omit, e. g., $\beta_{N-3}^{\textrm{Efn}}(\vec{r})$, and define the set $\{\bar{\beta}_{q}^{2}\}$ as
${\color[rgb]{0,0,0}\bar{\beta}_{q}^{2}(\vec{r})}=\begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...,N-4,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{1}1}(\vec{r}),&q=N-3,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{2}2}(\vec{r}),&q=N-2.\\ \end{cases}$ (79)
@@
$d_{n,11}(E)\!=\!\begin{cases}a_{n,11}^{as}(E)+g_{n,11}(E),&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}_{1}-1,\\ a_{n,11}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{1},...\,,\infty\end{cases}$ (80)
@@
$d_{n,12}(E)\!=\!\begin{cases}a_{n,12}^{as}(E)+g_{n,12}(E),\!\!&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}_{2\!}-1,\\ a_{n,12}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{2},...\,,\infty.\end{cases}$ (81)
@@
$d_{n,21}(E)\!=\!\begin{cases}a_{n,21}^{as}(E)+g_{n,21}(E),\!\!&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}_{1}-1,\\ a_{n,21}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{1},...\,,\infty\end{cases}$ (82)
@@
$d_{n,22}(E)=\begin{cases}a_{n,22}^{as}(E)+g_{n,22}(E),&n=0,...,\mathcal{\color[rgb]{0,0,0}N}_{2}-1,\\ a_{n,22}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{2},...,\infty.\end{cases}$ (83)
@@
$\!\!g_{n,ij}(E)$ $=\sum_{n^{\prime}=0}^{N-3}b_{n^{\prime}}^{i}(E)\,a_{n^{\prime}n}^{l_{j}},\!$ $i,j=1,2.$ (85)
In the case of the $np$ scattering, it is conventional to express the $S$ matrix in the Stapp parameterization Stapp where the phase shifts $\delta_{l_{i}}=\delta_{l_{i}}(E)$ are introduced in each of the partial waves and the mixing parameter $\epsilon=\epsilon(E)$. These parameters are related to the $S$ matrix as

### v2

@@ IV Two-Channel Scattering
We have shown that the adapted Efros method is consistent and leads to a convergence of single-channel scattering calculations using a limited basis space obtained by diagonalizing the truncated Hamiltonian $H^{tr}$. We now extend our application to a system with two open channels. In particular, we investigate $np$ scattering in the coupled $^{3}SD_{1}$ partial waves using the Daejeon16 $NN$ potential and eigenfunction SRFs.
[…]
@@
$H={\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j}-1}|\phi_{ni}\rangle V_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|+{\color[rgb]{0,0,0}\sum_{i=1,2}}\>\sum_{n,n^{\prime}=0}^{\infty}|\phi_{ni}\rangle T^{l_{i}}_{nn^{\prime}}\langle\phi_{n^{\prime}i}|\equiv{\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j}-1}|\phi_{ni}\rangle H^{tr}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|\\ +\sum_{i=1,2}\>\sum_{n,n^{\prime}=\mathcal{N}_{i}}^{\infty}|\phi_{ni\!}\rangle T^{l_{i}}_{nn^{\prime}}\langle\phi_{n^{\prime}i}\!|\!+\sum_{i=1,2}\!\big{[}|\phi_{\mathcal{N}_{i}-1,i\!}\rangle T^{l_{i}}_{\mathcal{N}_{i}-1,\mathcal{N}_{i}}\!\langle\phi_{\mathcal{N}_{i}i}\!|+|\phi_{\mathcal{N}_{i}i}\rangle T^{l_{i}}_{\mathcal{N}_{i},\mathcal{N}_{i}-1}\langle\phi_{\mathcal{N}_{i}-1,i}|\big{]}\!.$ (63)
We define the truncation of the potential energy matrix by the maximal oscillator quanta $N_{\max}$, i. e., ${2n+l_{i}}\leq N_{\max}$ and $\mathcal{N}_{i}=\frac{1}{2}(N_{\max}-l_{i})+1$, $i=1,2$.
@@
$\tilde{V}\equiv\sum_{i,j=1,2}\sum_{n=0}^{\mathcal{N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j}-1}|\phi_{ni}\rangle\tilde{V}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|=\sum_{i,j=1,2}\sum_{n=0}^{\mathcal{N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j}-1}|\phi_{ni}\rangle\sigma_{\mathcal{N}_{i}-1}^{n}V_{ni,n^{\prime}j}\sigma_{\mathcal{N}_{j}-1}^{n^{\prime}}\langle\phi_{n^{\prime}j}|.$ (64)
Smoothing also improves the convergence in the two-channel case and we present below results of calculations with the smoothed interactions. That means that we use the truncated Hamiltonian
$H^{tr}\equiv\tilde{H}^{tr}={\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j}-1}|\phi_{ni}\rangle H^{tr}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|$ (65)
@@
For the two-channel problem, it is convenient to solve Eq. (22) in the $S$-matrix representation. We define the set of SRFs $\{\beta_{q}\}$ as a set of N2N-2 eigenfunctions of $H^{tr}\!$,
@@
Note, $N=v+2$ is the number of unknowns in each channel and the total number of independent unknowns is ${\mathbb{\color[rgb]{0,0,0}N}}=2N-1\color[rgb]{0,0,0}=2v+3$}. We first solve the full set of equations for the $i=1$ entrance channel with the set $\{\bar{\beta}_{q}^{1}\}$ constituting a set of N2N-2 eigenfunctions of $H^{tr}$ plus two additional oscillator functions that are used exclusively for orthogonalization and guarantee coupling to the continuum but do not appear explicitly in Eq. (1):
${\color[rgb]{0,0,0}\bar{\beta}_{q}^{1}(\vec{r})}=\begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...,N-3,\\ \color[rgb]{0,0,0}\phi_{\mathcal{N}_{1}1}(\vec{r}),&q=N-2,\\ \color[rgb]{0,0,0}\phi_{\mathcal{N}_{2}2}(\vec{r}),&q=N-1.\end{cases}$ (68)
@@
$\beta_{q}^{\textrm{Efn}}(\vec{r})=\sum_{q^{\prime}=0}^{\mathcal{N}_{1}-1}a_{qq^{\prime}}^{l_{1}}\,{\color[rgb]{0,0,0}\phi_{q^{\prime}1}(\vec{r})}+\sum_{q^{\prime}=0}^{\mathcal{N}_{2}-1}a_{qq^{\prime}}^{l_{2}}\,\color[rgb]{0,0,0}\phi_{q^{\prime}2}(\vec{r}).$ (69)
Next, we solve the reduced set of equations for the i=2{i=2} entrance channel. Since there is one less unknown, namely, the off-diagonal $S$-matrix element $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$}, we use for the set $\{\bar{\beta}_{q}^{2}\}$ a subset of $\{\bar{\beta}_{q}^{1}\}$ that includes one less function. The choice of the function to be omitted for the orthogonalization in the $i=2$ channel is arbitrary. We can omit, e. g., $\beta_{N-3}^{\textrm{Efn}}(\vec{r})$, and define the set $\{\bar{\beta}_{q}^{2}\}$ as
[…]
@@
$d_{n,11}(E)\!=\!\begin{cases}a_{n,11}^{as}(E)+g_{n,11}(E),&n=0,...\,,\mathcal{N}_{1}-1,\\ a_{n,11}^{as}(E),&n=\mathcal{N}_{1},$$\infty\end{cases}$ (71)
@@
$d_{n,12}(E)\!=\!\begin{cases}a_{n,12}^{as}(E)+g_{n,12}(E),\!\!&n=0,...\,,\mathcal{N}_{2\!}-1,\\ a_{n,12}^{as}(E),&n=\mathcal{N}_{2},$$\infty.\end{cases}$ (72)
@@
$d_{n,21}(E)\!=\!\begin{cases}a_{n,21}^{as}(E)+g_{n,21}(E),\!\!&n=0,...\,,\mathcal{N}_{1}-1,\\ a_{n,21}^{as}(E),&n=\mathcal{N}_{1},$$\infty\end{cases}$ (73)
@@
$d_{n,22}(E)=\begin{cases}a_{n,22}^{as}(E)+g_{n,22}(E),&n=0,...,\mathcal{N}_{2}-1,\\ a_{n,22}^{as}(E),&n=\mathcal{N}_{2},$$\infty.\end{cases}$ (74)
@@
$\!\!g_{n,ij}(E)$ $=\sum_{n^{\prime}=0}^{v-1}b_{n^{\prime}}^{i}(E)\,a_{n^{\prime}n}^{l_{j}},\!$ $i,j=1,2.$ (76)
In the case of the $np$ scattering, it is conventional to express the $S$ matrix in the Stapp parametrization Stapp where the phase shifts $\delta_{l_{i}}=\delta_{l_{i}}(E)$ are introduced in each of the partial waves and the mixing parameter $\epsilon=\epsilon(E)$. These parameters are related to the $S$ matrix as